# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v2

@@ 4 EVALUATION
We first overview our experimental setup in Section 4.1 and then evaluate the performance of DP-ZO in Section 4.2. We find that DP-ZO provides a competitive privacy-utility trade-off for conservative privacy budgets across multiple datasets, model architectures and can scale to large models under conservative privacy budgets. We also compare DP-ZO to DP-SGD in Section 2 and show that DP-ZO achieves comparable performance to DP-SGD for the same model size. Furthermore, we show that DP-ZO achieves a non-trivial privacy-utility trade-off under pure $\varepsilon$-DP under a conservative privacy budget like $\varepsilon=4$ on large language models. In Section 4.3 we first ablates DP-ZO across different model architectures. We then measure the empirical privacy loss and computation efficiency of DP-ZO. We also characterize DP-ZO under different few-shot settings and different noise mechanisms for $(\varepsilon,\delta)$-DP.
@@ 4.1 Experimental Setup
We report the metric of interest (F1 score or accuracy) and standard deviation averaged across 5 independent runs with different random seeds. We detail the full hyperparameter searches in Appendix D.
@@
Datasets. We mainly consider three different benchmark NLP tasks: SQuAD (Rajpurkar et al., 2016) and DROP (Dua et al., 2019) for text generation, and SST2 (Socher et al., 2013) for text classification. Although all these datasets have very different dataset sizes, we consider the few-shot setting for all these datasets where we sample $1000$ points for each dataset. Fine-tuning LLMs with $O(n=1000)$ samples is a standard setting in the NLP community (Gao et al., 2021; Malladi et al., 2023) because we are generally interested in the few-shot abilities of LLMs (Brown et al., 2020). This represents a departure from prior works that privately finetune LLMs; Yu et al. (2022); Li et al. (2022b); Yu et al. (2021b) use the entire training dataset of SST2 that has about 65,000 examples. It is well known that the privacy-utility tradeoff improves greatly with more data (Tramèr and Boneh, 2021). It is straightforward to see that our setting with datasets of the size $n=1000$ with $\delta=10^{-5}$ is simultaneously more challenging and more aligned with real-world usecases than previous works in DP finetuning of LLMs. Despite the increased difficulty of our few-shot setting as compared to prior work, our results validate that DP-ZO realizes a strong privacy-utility trade-off. We also ablates the training sample size from the few-shot to the full training set by conducting experiments on the QNLI (Wang et al., 2019) dataset to be consistent with previous works (Li et al., 2022b; Yu et al., 2022) for a fair comparison.
@@
Models. We present our main results (Table 1) using a pretrained OPT-13B (Zhang et al., 2022) model that is finetuned with LoRA (Hu et al., 2022); that is, we update $<1\%$ of the total parameters. We include a range of ablation studies, including varying the model size among the OPT series, model architectures including Mistral-7B-v1 (Jiang et al., 2023) and amount of parameters to be updated, after we present the main results. We also include one experiments for QNLI on Roberta-base (Liu et al., 2019) to be consistent with previous works (Li et al., 2022b; Yu et al., 2022) for a fair comparison.
@@
Privacy budgets. We consider various privacy levels with $\varepsilon=[0.5,1,4]$ and fix $\delta=10^{-5}$ for $(\varepsilon,\delta)$-DP. We include the zero-shot $\varepsilon=0$ baseline that does not incur any privacy loss because we evaluate the pretrained model directly without finetuning on private data. We also include the non-private $\varepsilon=\infty$ baseline that is trained without any DP guarantee. That is, we iterate over the shuffled dataset instead of doing Poisson sampling (replacing line 4), do not clip the step size (skipping line 11) and set $\sigma=0$ (in line 13). We make these modifications because Poisson sampling and clipping are known to degrade performance, and we want to compare to the strongest possible nonprivate baseline.
@@ 4.2 Main Results
DP-ZO scales to large models. In Table 2 we show that DP-ZO continues improving as the model size increases from 1.3B to 66B. Due to space constraints, we provide the non-private ($\epsilon=\infty$) performance of all models and methods in Appendix E. Table 2 shows an promising insight: as the model size and nonprivate performance increase, the gap in performance between private and nonprivate models shrinks. Specifically, the gap for OPT-1.3B is $5.68\%$ ($80.97\%$ at $\varepsilon=\infty$ reduced to $75.29\%$ under $\varepsilon=1$). But this gap shrinks to just $3.37\%$ for OPT-66B, where the private performance at $\epsilon=1$ is $84.12\%$ compared to $87.49\%$ non-privately. Our findings suggest that DP-ZO scales to large models not only because it is compatible with existing pipeline without much additional engineering effort but also because the utility drop due to privacy is smaller as the model size increases.
@@
Comparison with DP-SGD. We compare DP-ZO to differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016) which has been applied to fine-tune LLMs with full parameter fine-tuning (Li et al., 2022b) and with LoRA (Yu et al., 2022; He et al., 2023). Recall that DP-ZO is compatible out-of-the-box with mixed precision training and GPU parallelism, enabling us to fine-tune OPT-66B. As we discuss in  Section 5, it is significantly more challenging to integrate DP-SGD with these techniques, and furthermore, DP-SGD requires more memory than DP-ZO to store activatiosn and compute per-sample gradients (see Section 4.3). As a direct result, DP-SGD cannot directly scale past 2.7B with full fine-tuning or 13B with LoRA without additional implementation effort for multi-GPU training, while DP-ZO can scale seamlessly to larger models. In Table 2 we present comparisons between DP-ZO and DP-SGD with full parameter finetuning and LoRA. With the same model size, DP-ZO achevies comparable performance as DP-SGD as by LoRA finetuning, i.e., both DP-ZO and DP-SGD achieves $82\%$ on OPT-13B models. The best performance by DP-ZO is $84.12\%$ by OPT-66B finetuned with LoRA. This is $\approx 2\%$ better than the best performance of DP-SGD in Table 2 that is $82.14\%$ by OPT-13B with LoRA.
@@
In Table 3, DP-ZO (Laplace) can significantly improve upon $\varepsilon=0$. Given a budget $\varepsilon=4$, which some prior work has considered reasonable (Ponomareva et al., 2023), DP-ZO (Laplace) can obtain $73.52\%$ on SQuAD. When increasing $\varepsilon=4$ to $\varepsilon=15$, DP-ZO (Laplace) can obtain $78.82\%$ on SQuAD. Note that the $l_{1}$ sensitivity required for Laplace mechanism makes it hard to DP-SGD to achieve comparable performance as DP-ZO because the gradients in DP-SGD have high dimension. Table 3 shows that DP-SGD with $l_{1}$ norm clipping and Laplace noise only achieves $47.25\%$ for reasonable privacy budgets with $\varepsilon$ ranging from $4$ to $15$, that is only marginal improvement upon the zero-shot performance. Even when relaxing the privacy budget to near-vacuous guarantees such as $\varepsilon=10450$, DP-SGD (Laplace) still achieves worse performance compared to DP-ZO due to the Laplace noise added in high-dimension gradients.
@@ 4.3 Analysis
In this section, we provide ablation studies on the feasibility of DP-ZO across different model architectures, empirical privacy analysis to measure how private is DP-ZO, memory efficiency of DP-ZO, the amount of training data that we sample in DP-ZO, and the choice of DP mechanism in DP-ZO.
DP-ZO provides a strong privacy-utility trade-off across different model architectures. Table 1 and Table 2 show that DP-ZO achieves the comparable performance as DP-SGD on various OPT models sizes. We now run experiments on SQuAD with Mistral-7B-v1 model (Jiang et al., 2023) in Table 4 and include the results of OPT-6.7B and OPT-13B for the ease of comparison. DP-ZO and DP-SGD both achieve comparable performance at $\varepsilon=1$, that are around $89\%$. Moreover, even Mistral-7B-v1 has similar model parameters as OPT-6.7B and much fewer parameters than OPT-13B, DP-ZO achieves better performance by Mistral-7B-v1 than OPT-13B. This indicates that with the development of more advanced models, the power of DP-ZO will be further unlocked.
Table 4: Ablating DP-ZO across different model architectures. The $\varepsilon=0$ baseline for Mistral-7B-v1 is 68.37.
@@
DP-ZO is memory efficient. We omitted several results of DP-SGD in Section 4.2 due to excessive memory consumption of DP-SGD that leads to out-of-memory (OOM) on a single A100 80G GPU. We now provide a more fine-grained memory analysis for a better understanding of DP-ZO and DP-SGD.
@@
As discussed in (Li et al., 2022b; De et al., 2022), small batch size will incur sub-optimal performance of DP-SGD and therefore large batch size is preferred, we consider gradient accumulation in memory analysis to enable large batch size. We consider full precision both for DP-SGD and DP-ZO for fair comparison. We present the results of such memory consumption for different sequence lengths on OPT-2.7B in Table 5.
Table 5: Comparison of memory consumption (GB) of DP-ZO and DP-SGD varying differnet sequence length on OPT-2.7B. Full-precision, no gradient check-pointing. Batch size=2, gradient accumulation steps=2. OOM indicates out-of-memory on a single A100 80G GPU.
@@
seq_len=128 $51.3$ $32.9$ $11.4$ $12.2$ $11.6$ $10.3$
seq_len=512 $51.4$ $39.6$ $18.1$ $21.3$ $11.7$ $11.1$
seq_len=2048 OOM OOM OOM OOM $15.6$ $15.6$
Table 5 shows that the naive implementation of DP-SGD incurs around 50GB for sequence length equal to 128 and incurs out-of-memory (OOM) issue when increasing sequence length to 2048. Ghost clipping can help reduce the memory consumption by removing per-sample gradient computation, but still needs to accumulate gradients and consumes memory more than 30GB. Parameter efficient fine-tuning methods including DP-SGD-LoRA, DP-SGD-BitFiT can largely reduce the memory cost for gradient accumulation with fewer parameters in gradients. However, as the input sequence length increases, the activation saved in forward-pass for gradients computation significantly increases and still causes OOM error for sequence length equals to 2048. Note that such long sequence input exists in practical scenarios such as digesting from a long documents and recent foundation models put efforts to support longer context (Gemini-Team et al., 2023). In contrast, DP-ZO does not need to store gradients nor store additional activation for gradients, therefore can reduce the memory consumption to be less than 16GB even when sequence length is 2048. In fact, DP-ZO incurs nearly no additional memory cost than ZO (Malladi et al., 2023).
@@
We now consider a more restrictive memory setting such as on-device mobile setting, i.e., 8GB as a memory limit (Gim and Ko, 2022; Guo et al., 2024). As discussed in Malladi et al. (2023), gradient checkpointing can help reduce the activation memory consumption. We use gradient checkpointing to reduce the memory consumption of activation and half-precision to reduce the memory consumption of weights and gradients. We measure with the default implementation of activation checkpointing where we checkpoint every block of the model. We consider batch size=1 and accumulate gradients in 2 steps to enable training in large batch size. We present the results of such memory consumption for different sequence length on OPT-2.7B model in Figure 5. While gradient checkpointing largely reduces the memory consumption in DP-SGD, DP-SGD-full is still beyond the 8GB memory limit due to gradient accumulation. Similarly, with gradient checkpointing, DP-SGD-LoRA still exceeds the 8GB limit when increasing sequence to 2048. In contrast, DP-ZO incurs just 6.2GB even when input sequence length is 2048.
@@
How private is DP-ZO and ZO: an empirical privacy analysis. We discussed earlier in Section 2.2 that the single scalar information from Zeroth-order Optimization will leak private information and therefore motivate our design of DP-ZO. We now validate our design by conducting empirical privacy evaluation through membership inference attacks (MIA) (Shokri et al., 2017) to understand the privacy implication of DP-ZO. We use the state-of-the-art privacy auditing empirical privacy evaluation method in Panda et al. (2024b) for empirical privacy evaluation. Similar to Panda et al. (2024b), we construct synthetic canaries by creating one new token for each canary to the vocabulary to increase the information from canaries for better auditing. Following Panda et al. (2024b); Mireshghallah et al. (2022), we use MIA Area under the ROC Curve (AUC-ROC) for empirical privacy evaluation. We report the results for DP-ZO and DP-SGD in Table 6 for different $\varepsilon$s including $[0.5,1,4,10,\infty]$.
@@
As observed in Panda et al. (2024b), such privacy attack is very successful for DP-SGD when no noise added, and DP-SGD can effectively reduce the attack AUC to 54.8 that is closed to random guess at $\varepsilon=0.5$. For DP-ZO, when no noise added, the MIA AUC is 71.4 that is lower than DP-SGD($\varepsilon=\infty$), however much higher than the random guess baseline 50. Interestingly, in this empirical privacy case study, the privacy leakage of ZO is similar to the privacy leakage of DP-SGD at $\varepsilon=10$. DP-ZO is motivated to get a formal privacy guarantee by differential privacy. DP-ZO can reduce the attack AUC to around random guess at $\varepsilon=[0.5,1,4]$. To the best of our knowledge, this is the first experimental result that measures the privacy leakage in ZO and DP-ZO. This result shows the necessity of DP-ZO to redcue MIA cloase to random guess, and also raises an open problem about the inherent privacy property of zeroth-order optimization.
Table 6: Membership Inference Atttack AUC-ROC for DP-ZO and DP-SGD.
@@
Characterizing the effect of data size. Although it is known that private learning requires more data than non-private learning (Bassily et al., 2014), prior work has not characterized this improvement for fine-tuning language models. In Table 8 and Figure 6 we first vary the number of training samples $n$ around the $n=1000$ setting in the main results while keeping $\delta=10^{-5}$ fixed for all choices of $n$. Table 8 shows that DP-ZO can achieve nontrivial performance in few-shot settings under conservative privacy guarantees. Furthermore, we find that while increasing the amount of training data by 10×10$\times$10 × barely increases non-private performance, it increases private performance by $\approx 6\%$ ($n=500$ vs. $n=5000$). Similarly, acquiring more data enhances privacy amplification and reduces the amount of noise we need to add to achieve a target $\varepsilon$-DP + guarantee. In Table 8 we find that increasing the number of training examples from $1000$ to $5000$ improves performance at $\varepsilon=4$ from $73.52\%$ to $79.89\%$, although the improvement of non-private performance at $\epsilon=\infty$ by increasing training samples from $1000$ to $5000$ is insignificant.
@@
While non-private few-shot learning can succeed by just memorizing the training data, Figure 6 indicates that the convergence rate for different shots for private few-shot learning is different. With the proliferation of pretrained models, we anticipate that privately fine-tuning downstream tasks in the few-shot setting will be more aligned with real-world use cases.
Besides the few-shot setting, we now scale the size of training data size up to more than 100k samples and compare the performance of DP-SGD and DP-ZO. For the ease of comparison, we follow Li et al. (2022b); Yu et al. (2022) and conduct experiments for DP-ZO and DP-SGD on QNLI with RoBERTa-base with a range of examples at $\varepsilon=3$ and report result in Table 10. Similar to previous observation, Table 10 shows that at a small data regime, i.e., samples=1000, DP-ZO achieves comparable performance as DP-SGD. DP-ZO can also improve its performance within more samples. Compared to DP-SGD, we notice that when increasing the data size, there is a utility gap between DP-ZO and DP-SGD. This limitation shows an open problem that how to improve data efficiency in zeroth-order optimization (Zhang et al., 2024c; Zhao et al., 2024) and DP-ZO, and we leave this data efficiency improvement in ZO as future work.
Table 7: Ablation of DP-ZO (Gaussian) for different $n$ training samples on SQuAD dataset. ($1,10^{-5}$)-DP. OPT-13B with LoRA finetuning.
@@
Table 8: Pure $\varepsilon$-DP by DP-ZO (Laplace) at $\varepsilon=4$, SQuAD with different training samples. OPT-13B with LoRA fine-tuning. The $\varepsilon=\infty$ by ZO is $86.85\%$ and $86.92\%$ for 1000 and 5000 samples respectively. The $\varepsilon=0$ baseline is $46.2\%$.
@@
Table 9: Comparison of DP-ZO and DP-SGD on different samples on QNLI dataset with Roberta-base model.
1000 5000 10000 50000 104743
@@
Table 10: DP-ZO with different DP mechanism. SQuAD with 1000 training samples. $\delta=10^{-5}$. G is for Gaussian and L is for Laplace.

### v3

@@ 4 EVALUATION
We first overview our experimental setup in Section 4.1 and then evaluate the performance of DP-ZO in Section 4.2. We find that DP-ZO provides a competitive privacy-utility trade-off for conservative privacy budgets across multiple datasets, model architectures and can scale to large models under conservative privacy budgets. We also compare DP-ZO to DP-SGD in Section 2 and show that DP-ZO achieves comparable performance to DP-SGD for the same model size. Furthermore, we show that DP-ZO achieves a non-trivial privacy-utility trade-off under pure $\varepsilon$-DP under a conservative privacy budget like $\varepsilon=4$ on large language models. In Section 4.3 we first provides results of DP-ZO across different model architectures. We then measure the empirical privacy loss and computation efficiency of DP-ZO. We also characterize DP-ZO under different few-shot settings and different noise mechanisms for $(\varepsilon,\delta)$-DP.
@@ 4.1 Experimental Setup
We report the metric of interest (F1 score or accuracy) and standard deviation averaged across 5 independent runs with different random seeds. We detail the full hyper-parameter searches and computation cost in Section D.2.
@@
Datasets. Following Malladi et al. (2023), we mainly consider three different benchmark NLP tasks: SQuAD (Rajpurkar et al., 2016) and DROP (Dua et al., 2019) for text generation, and SST2 (Socher et al., 2013) for text classification. We use F1 for text generation and accuracy for text classification as evaluation metric (we include a detailed description of the metric in Section D.1). Although all these datasets have very different dataset sizes, we consider the few-shot setting for all these datasets where we sample $1000$ points for each dataset. Fine-tuning LLMs with $O(n=1000)$ samples is a standard setting in the NLP community (Gao et al., 2021; Malladi et al., 2023) because we are generally interested in the few-shot abilities of LLMs (Brown et al., 2020). This represents a departure from prior works that privately finetune LLMs; Yu et al. (2022); Li et al. (2022b); Yu et al. (2021b) use the entire training dataset of SST2 that has about 65,000 examples. It is well known that the privacy-utility trade-off improves greatly with more data (Tramèr & Boneh, 2021). It is straightforward to see that our setting with datasets of the size $n=1000$ with $\delta=10^{-5}$ is simultaneously more challenging and more aligned with real-world usecases than previous works in DP finetuning of LLMs. Despite the increased difficulty of our few-shot setting as compared to prior work, our results validate that DP-ZO realizes a strong privacy-utility trade-off. We also varies the training sample size from the few-shot to the full training set by conducting experiments on the QNLI (Wang et al., 2019) dataset to be consistent with previous works (Li et al., 2022b; Yu et al., 2022) for a fair comparison.
@@
Models. We present our main results (Table 1) using a pretrained OPT-13B (Zhang et al., 2022) model that is finetuned with LoRA (Hu et al., 2022); that is, we update $<1\%$ of the total parameters. We include a range of analysis, including varying the model size among the OPT series, model architectures including Mistral-7B-v1 (Jiang et al., 2023) and amount of parameters to be updated, after we present the main results. We also include one experiments for QNLI on RoBERTa-base (Liu et al., 2019) to be consistent with previous works (Li et al., 2022b; Yu et al., 2022) for a fair comparison.
@@
Privacy budgets. We consider various privacy levels with $\varepsilon=[0.5,1,4]$ and fix $\delta=10^{-5}$ for $(\varepsilon,\delta)$-DP. We include the zero-shot $\varepsilon=0$ that does not incur any privacy loss because we evaluate the pretrained model directly without finetuning on private data. We also include the non-private $\varepsilon=\infty$ baseline that is trained without any DP guarantee. That is, we iterate over the shuffled dataset instead of doing Poisson sampling (replacing line 3), do not clip the step size (skipping line 10) and set $\sigma=0$ (in line 13). We make these modifications because Poisson sampling and small threshold like $C=0.1$ for per-example clipping makes gradient estimator biased and loss convergence issues in $\sigma=0$ for DP-SGD (Andrew et al., 2021; Chen et al., 2020; De et al., 2022; Bu et al., 2023b), and we want to compare to the strongest possible nonprivate baseline.
@@ 4.2 Main Results
DP-ZO scales to large models. In Table 2 we show that DP-ZO continues improving as the model size increases from 1.3B to 66B. Due to space constraints, we provide the non-private ($\varepsilon=\infty$) performance of all models and methods in Appendix E. Table 2 shows an promising insight: as the model size and non-private performance increase, the gap in performance between private and non-private models shrinks. Specifically, the gap for OPT-1.3B is $5.68\%$ ($80.97\%$ at $\varepsilon=\infty$ reduced to $75.29\%$ under $\varepsilon=1$). But this gap shrinks to just $3.37\%$ for OPT-66B, where the private performance at $\epsilon=1$ is $84.12\%$ compared to $87.49\%$ non-privately. Our findings suggest that DP-ZO scales to large models not only because it is compatible with existing pipeline without much additional engineering effort but also because the utility drop due to privacy is smaller as the model size increases.
@@
Comparison with DP-SGD. We compare DP-ZO to differentially private stochastic gradient descent (DP-SGD) (Abadi et al., 2016) which has been applied to fine-tune LLMs with full parameter fine-tuning (Li et al., 2022b) and with LoRA (Yu et al., 2022; He et al., 2023). Recall that DP-ZO is compatible out-of-the-box with mixed precision training and GPU parallelism, enabling us to fine-tune OPT-66B. As we discuss in  Section 5, it is significantly more challenging to integrate DP-SGD with these techniques, and furthermore, DP-SGD requires more memory than DP-ZO to store activations and compute per-sample gradients (see Section 4.3). As a direct result, DP-SGD cannot directly scale past 2.7B with full fine-tuning or 13B with LoRA without additional implementation effort for multi-GPU training, while DP-ZO can scale seamlessly to larger models. In Table 2 we present comparisons between DP-ZO and DP-SGD with full parameter finetuning and LoRA. With the same model size, DP-ZO achieves comparable performance as DP-SGD as by LoRA finetuning, i.e., both DP-ZO and DP-SGD achieves $82\%$ on OPT-13B models. The best performance by DP-ZO is $84.12\%$ by OPT-66B finetuned with LoRA. This is $\approx 2\%$ better than the best performance of DP-SGD in Table 2 that is $82.14\%$ by OPT-13B with LoRA.
@@
In Table 3, DP-ZO (Laplace) can significantly improve upon $\varepsilon=0$. Given a budget $\varepsilon=4$, which some prior work has considered reasonable (Ponomareva et al., 2023), DP-ZO (Laplace) can obtain $73.52\%$ on SQuAD. When increasing $\varepsilon=4$ to $\varepsilon=15$, DP-ZO (Laplace) can obtain $78.82\%$ on SQuAD. Note that the $l_{1}$ sensitivity required for Laplace mechanism makes it hard for DP-SGD to achieve comparable performance as DP-ZO because the gradients in DP-SGD have high dimension. Table 3 shows that DP-SGD with $l_{1}$ norm clipping and Laplace noise only achieves $47.25\%$ for reasonable privacy budgets with $\varepsilon$ ranging from $4$ to $15$, that is only marginal improvement upon the zero-shot performance. Even when relaxing the privacy budget to near-vacuous guarantees such as $\varepsilon=10450$, DP-SGD (Laplace) still achieves worse performance compared to DP-ZO due to the Laplace noise added in high-dimension gradients.
@@ 4.3 Analysis
In this section, we provide experiments for DP-ZO across different model architectures, empirical privacy analysis to measure how private is DP-ZO, memory efficiency of DP-ZO, the amount of training data that we sample in DP-ZO, and the choice of DP mechanism in DP-ZO.
DP-ZO provides a strong privacy-utility trade-off across different model architectures. Table 1 and Table 2 show that DP-ZO achieves the comparable performance as DP-SGD on various OPT models sizes. We now run experiments on SQuAD with Mistral-7B-v1 model (Jiang et al., 2023) in Table 4 and include the results of OPT-6.7B and OPT-13B for the ease of comparison. On Mistral-7B-v1, DP-ZO and DP-SGD both achieve comparable performance at $\varepsilon=1$, that are around $89\%$. Moreover, even Mistral-7B-v1 has similar model parameters as OPT-6.7B and much fewer parameters than OPT-13B, DP-ZO achieves better performance by Mistral-7B-v1 than OPT-13B. This indicates that with the development of more advanced models, the power of DP-ZO will be further unlocked.
Table 4: The results of DP-ZO on different model architectures. The $\varepsilon=0$ baseline for Mistral-7B-v1 is 68.37.
@@
DP-ZO is memory efficient. We omitted several results of DP-SGD in Section 4.2 due to excessive memory consumption of DP-SGD that leads to out-of-memory (OOM) issue on a single A100 80G GPU. We now provide a more fine-grained memory analysis for a better understanding of DP-ZO and DP-SGD.
@@
As discussed in Li et al. (2022b); De et al. (2022), small batch size will incur sub-optimal performance of DP-SGD and therefore large batch size is preferred, we consider gradient accumulation in memory analysis to enable large batch size. We consider full precision both for DP-SGD and DP-ZO for fair comparison. We present the results of such memory consumption for different sequence lengths on OPT-2.7B in Table 5.
Table 5: Comparison of memory consumption (GB) of DP-ZO and DP-SGD varying different sequence length $L$ on OPT-2.7B. Full-precision, no gradient check-pointing. Batch size=2, gradient accumulation steps=2. OOM indicates out-of-memory on a single A100 80G GPU.
@@
$L=128$ $51.3$ $32.9$ $11.4$ $12.2$ $11.6$ $10.3$
$L=512$ $51.4$ $39.6$ $18.1$ $21.3$ $11.7$ $11.1$
$L=2048$ OOM OOM OOM OOM $15.6$ $15.6$
Table 5 shows that the naive implementation of DP-SGD incurs around 50GB for sequence length $L=128$ and incurs out-of-memory (OOM) issue when increasing sequence length to $L=2048$. Ghost clipping can help reduce the memory consumption by removing per-sample gradient computation, but still needs to accumulate gradients and consumes memory more than 30GB. Parameter efficient fine-tuning methods including DP-SGD-LoRA, DP-SGD-BitFiT can largely reduce the memory cost for gradient accumulation with fewer parameters in gradients. However, as the input sequence length $L$ increases, the activation saved in forward-pass for gradients computation significantly increases and inputs with sequence length $L=2048$ still cause OOM error for parameter efficient fine-tuning methods. Note that such long sequence input exists in practical scenarios such as digesting from a long documents and recent foundation models put efforts to support longer context (Gemini-Team et al., 2023). In contrast, DP-ZO does not need to store gradients nor store additional activation for gradients, therefore can reduce the memory consumption to be less than 16GB even when sequence length is 2048. In fact, DP-ZO only incurs a negligible additional memory cost than ZO (Malladi et al., 2023) for the per-sample loss clipping and noise addition.
@@
We now consider a more restrictive memory setting such as on-device mobile setting, i.e., 8GB as a memory limit (Gim & Ko, 2022; Guo et al., 2024). As discussed in Malladi et al. (2023), gradient checkpointing can help reduce the activation memory consumption. We use gradient checkpointing to reduce the memory consumption of activation and half-precision to reduce the memory consumption of weights and gradients. We measure the memory cost with the default implementation of activation checkpointing where we checkpoint every block of the model. We consider batch size=1 and accumulate gradients in 2 steps to enable training in large batch size. We present the results of the memory consumption for different sequence lengths on OPT-2.7B model in Figure 5. While gradient checkpointing significantly reduces the memory consumption in DP-SGD, DP-SGD-full is still beyond the 8GB memory limit due to gradient accumulation. Similarly, with gradient checkpointing, DP-SGD-LoRA still exceeds the 8GB limit when increasing sequence length to 2048. In contrast, DP-ZO incurs just 6.2GB even when input sequence length is 2048.
@@
How private is DP-ZO and ZO: an empirical privacy analysis. We discussed earlier in Section 2.2 that the single scalar information from Zeroth-order Optimization will leak private information and therefore motivate our design of DP-ZO. We now validate our design by conducting empirical privacy evaluation through membership inference attacks (MIA) (Shokri et al., 2017) to understand the privacy implication of DP-ZO. We use the state-of-the-art privacy auditing method in Panda et al. (2024b) for empirical privacy evaluation. Similar to Panda et al. (2024b), we construct synthetic canaries by creating one new token for each canary to the vocabulary to increase the information from canaries for better auditing. Following Panda et al. (2024b); Mireshghallah et al. (2022), we use MIA Area under the ROC Curve (AUC-ROC) for empirical privacy evaluation. We report the results for DP-ZO and DP-SGD in Table 6 for different $\varepsilon$s including $[0.5,1,4,10,\infty]$.
@@
As observed in Panda et al. (2024b), this privacy attack method is very successful for DP-SGD when no noise added, and DP-SGD can effectively reduce the attack AUC to 54.8 that is closed to random guess at $\varepsilon=0.5$. For DP-ZO, when no noise added, the MIA AUC is 71.4 that is lower than DP-SGD($\varepsilon=\infty$), however much higher than the random guess baseline 50. Interestingly, in this empirical privacy case study, the privacy leakage of ZO is similar to the privacy leakage of DP-SGD at $\varepsilon=10$. DP-ZO is motivated to get a formal privacy guarantee by differential privacy. DP-ZO can reduce the attack AUC to around random guess at $\varepsilon=[0.5,1,4]$. To the best of our knowledge, this is the first experimental result that measures the privacy leakage in ZO and DP-ZO. This result shows the necessity of DP-ZO to reduce MIA close to random guess, and also raises an open problem about the inherent privacy property of zeroth-order optimization.
Table 6: Membership Inference Attack AUC-ROC for DP-ZO and DP-SGD.
@@
Characterizing the effect of data size. Although it is known that private learning requires more data than non-private learning (Bassily et al., 2014), prior work has not characterized this improvement for fine-tuning language models. In Table 8 and Figure 6 we first vary the number of training samples $n$ around the $n=1000$ setting in the main results while keeping $\delta=10^{-5}$ fixed for all choices of $n$. Table 8 shows that DP-ZO can achieve nontrivial performance in few-shot settings under conservative privacy guarantees. Furthermore, we find that while increasing the amount of training data by 10×10$\times$10 × barely increases non-private performance, it increases private performance by $\approx 6\%$ ($n=500$ vs. $n=5000$). Similarly, for $\varepsilon$-DP, acquiring more data enhances privacy amplification and reduces the amount of noise we need to add to achieve a target DP $\varepsilon$ guarantee. In Table 8 we find that increasing the number of training examples from $1000$ to $5000$ improves performance at $\varepsilon=4$ from $73.52\%$ to $79.89\%$, although the improvement of non-private performance at $\epsilon=\infty$ by increasing training samples from $1000$ to $5000$ is insignificant.
While non-private few-shot learning can succeed by just memorizing the training data, Figure 6 indicates that the convergence rate for different shots for private few-shot learning is different. With the proliferation of pretrained models, we anticipate that privately fine-tuning downstream tasks in the few-shot setting will be more aligned with real-world use cases (Gao et al., 2021).
Table 7: The effect of different $n$ training samples for DP-ZO (Gaussian) fon SQuAD dataset. ($1,10^{-5}$)-DP. OPT-13B with LoRA finetuning.
@@
Table 8: $\varepsilon$-DP ($\varepsilon=4$) by DP-ZO (Laplace) on different $n$ in SQuAD on OPT-13B (LoRA). . $\varepsilon=\infty$ by ZO is $86.85\%$ for $n=1000$ and $86.92\%$ for $n=5000$. $\varepsilon=0$ is $46.2\%$.
@@
Table 9: Comparison of DP-ZO and DP-SGD on different $n$ in QNLI on RoBERTa-base model. $(3,4.7\times 10^{-6})$-DP.
$n$-shot 1000 5000 10000 50000 104743
@@
Table 10: DP-ZO with different DP mechanism. SQuAD with 1000 training samples. $\delta=10^{-5}$.